### v1

# CyC3D: Fine-grained Controllable 3D Generation via Cycle Consistency Regularization

Abstract
Despite the remarkable progress of 3D generation, achieving controllability, i.e., ensuring consistency between generated 3D content and input conditions like edge and depth, remains a significant challenge. Existing methods often struggle to maintain accurate alignment, leading to noticeable discrepancies. To address this issue, we propose CyC3D, a new framework that enhances controllable 3D generation by explicitly encouraging cyclic consistency between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls. Specifically, we employ an efficient feed-forward backbone that can generate a 3D object from an input condition and a text prompt. Given an initial viewpoint and a control signal, a novel view is rendered from the generated 3D content, from which the extracted condition is used to regenerate the 3D content. This re-generated output is then rendered back to the initial viewpoint, followed by another round of control signal extraction, forming a cyclic process with two consistency constraints. View consistency ensures coherence between the two generated 3D objects, measured by semantic similarity to accommodate generative diversity. Condition consistency aligns the final extracted signal with the original input control, preserving structural or geometric details throughout the process. Extensive experiments on popular benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
1 Introduction
Figure 1: Visualization of input and reconstructed conditions. We feed the input condition to the generation model first and use the generated 3D contents to render the image at the input viewpoint to extract the reconstructed conditions. The difference between the input and the reconstructed conditions can reveal whether the fine-grained controls are reserved in the results.
Figure 2: Brief overview of our proposed CyC3D. We first feed the condition on the original reference view to the generation model and extract the conditions from the rendered results on randomly sampled novel views. The extracted conditions are then fed to the generation model again and rendered back to the original reference view. The extracted condition of the rendered results should be consistent with the input condition on the reference view.
The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR [2], digital game design [32], and filmmaking [30]. The mainstream approaches can be broadly categorized into two groups: optimization-based methods and feed-forward methods. Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) [20] to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models [24]. In contrast, feed-forward 3D generation approaches [19, 28, 11, 12, 36, 29, 5] generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
While significant progress has been made in 3D generation research, a critical yet under-explored challenge remains—the controllability of 3D generation. Controllable generation seeks to integrate precise control signals, such as edges, sketches, depth maps, surface normals, and text prompts, as conditioning inputs. By incorporating these structured constraints, controllable 3D generation enhances the flexibility and reliability of the synthesis process, enabling more precise manipulation of the generated content.
To advance research in controllable 3D generation, MVControl [15] extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model [25]. It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model [27] to generate the final 3D representation. In contrast, ControLRM [31] aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) with a conditional encoder, enabling more precise conditioning. While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., achieving precise and fine-grained conditional control. As illustrated in Fig. 1, the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
To address this challenge, we propose CyC3D, a novel and general training framework that enhances the controllability of 3D generation. Our approach extends a feed-forward 3D generation network into a cyclic paradigm, enabling more precise and consistent adherence to input control signals, as illustrated in Fig. 2. Specifically, given an input control condition on a reference view, the generated 3D content is first rendered from randomly sampled new viewpoints. From these rendered images, the corresponding control conditions are extracted and fed back into the generation model as inputs, ensuring that the model learns to maintain control consistency across views. Subsequently, the model regenerates the 3D content, which is then rendered back to the initial reference viewpoint, where the extracted control conditions should match the original input condition. To further refine controllability, we introduce two complementary consistency constraints: View Cycle Consistency, which ensures that the two generated contents have similar semantics in embedded space since they are generated with the same prompt and highly coupled conditions, and Condition Cycle Consistency, which enforces strict alignment between the extracted control signals from the final rendered images and the original input condition. We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets [31] for zero-shot evaluation. Experimental results demonstrate that our framework significantly enhances controllability across all benchmarks while preserving competitive generation quality compared to existing methods.
We highlight the contributions of this work as follows:
  • New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency (Fig. 2).
  • Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency. The former regularizes the consistency between input and reconstructed conditions, the latter enhances the semantic consistency of different stages.
  • Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
2 Related Work
Text-/Image-to-3D Generation. 3D generation can be broadly categorized into optimization-based and feed-forward approaches: (1) Optimization-based methods offer a viable alternative to traditional techniques by circumventing the need for extensive text-to-3D datasets. DreamFusion [20] exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets. Despite their effectiveness, these methods often encounter the Janus problem, which arises from inconsistencies between 2D supervision and 3D geometry. (2) Feed-forward models have emerged as a prominent approach for efficient 3D generation, leveraging extensive 3D datasets. Techniques such as LRM [12] and TripoSR [29] utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities. Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer [16] and MVDream [25].
Controllable 3D Generation. Achieving controllability in 3D generation remains a critical objective in the field. Techniques developed for controllable 2D generation have provided valuable insights and methods that are applicable to the more complex domain of controllable 3D generation. Inspired by ControlNet [35], MVControl [15] integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation. This method combines coarse Gaussian generation with SDS-based refinement but faces challenges related to model generalization and procedural complexity. ControLRM [31] addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation. Building on the advances of ControlNet++ [14] and ControLRM [35], we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
3 CyC3D
Figure 3: Overall framework of our proposed CyC3D. (a) summarize the procedure of our CyC3D. (b) shows the condition extractor utilized for different conditional controls (Sec. 3.2). (c) illustrates the condition cycle consistency regularization (Sec. 3.3). (d) introduces the proposed view cycle consistency regularization (Sec. 3.4).
As shown in Fig. 3, we present the overall framework of our proposed CyC3D. In Sec. 3.1, we first introduce the feed-forward backbone network for controllable 3D generation in our CyC3D. Then we introduce the condition extractor (Fig. 3(b)) for different conditional controls (e.g. edge, sketch, depth, normal) in Sec. 3.2. In Sec. 3.3, we introduce the design of condition cycle consistency regularization (Fig. 3(c)) to enhance the fined-grained controls. In Sec. 3.4, we further introduce the proposed view cycle consistency regularization (Fig. 3(d)) to stabilize the training and avoid overfitting on the input condition. Finally, we introduce the overall loss in Sec. 3.5.
3.1 Generation Backbone
For controllable 3D generation, our CyC3D can be combined with arbitrary feed-forward backbones (e.g. ControLRM [31], MVControl [15]). In default, we select ControLRM [31] as our backbone model. ControLRM contains a 2D condition extractor and a 3D triplane decoder. (1) The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs. This generator can utilize either a transformer or diffusion backbone to create initial 2D latent representations, which are subsequently transformed into 3D models. (2) The 3D triplane decoder is inherited from a pre-trained large reconstruction model [12]. The triplane transformer decoder employs a unique representation of 3D data by decomposing it into three orthogonal planes (X, Y, and Z). Each plane is processed using transformer layers, which utilize self-attention mechanisms to capture spatial relationships and dependencies within the data.
3.2 Condition Extractor
To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor in Fig. 2 should be differentiable. The design of the condition extractor for each condition control is shown as follows:
Canny Condition. As shown in the figure, we extract edge maps using the Canny algorithm [4]. Following [14], a differentiable Canny extractor can implemented based on Kornia [23] to extract edge maps.
Sketch Condition. Following ControlNet [35], we utilize a pre-trained CNN network to extract sketch conditions from images. Since the CNN network is differentiable, the gradients can be passed backward through the network and back-propagated to the remaining modules.
Depth Condition. In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything [33]) to extract the depth condition from rendered images. The whole process is differentiable in the cycle consistency pipeline.
Normal Condition. In practice, we find that the existing normal estimation models [1] utilized in ControlNet [35] tends to predict over-smoothed normal maps. It may result in a large gap between the exact 3D shape of the rendered multi-view images and the estimated normal maps. Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to normal map via a differentiable module [13]. The great zero-shot generalization ability of large depth-estimation models like Depth-Anything [33] can effectively capture the 3D prior from the image and thus generate accurate normal map.
3.3 Condition Cycle Consistency Regularization
Denote the input edge condition as $C_{e}\in\mathbb{R}^{H\times W}$, the sketch condition as $C_{s}\in\mathbb{R}^{H\times W}$, the depth condition as $C_{d}\in\mathbb{R}^{H\times W}$, and the normal condition as $C_{n}\in\mathbb{R}^{H\times W\times 3}$. The input visual condition is noted as $C_{v}\in\{C_{e},C_{s},C_{d},C_{n}\}$. Given the text prompt $C_{t}$ and the input visual condition as $C_{v}$, the generation model outputs the triplane representation $P\in\mathbb{R}^{3\times C_{p}\times H_{p}\times W_{p}}$ corresponding to the input conditions.
$\small P_{i}=f_{\text{gen}}(C_{v},C_{t},T_{i}),$ (1)
where $T_{i}$ is the camera pose of the input condition, which is usually assumed to be an identity matrix after normalizing all the camera extrinsics to a frontal view. $f_{\text{gen}}$ is the backbone generation model for controllable 3D generation.
Given a 3D point $x\in\mathbb{R}^{3}$, the color and density filed $(c,\sigma)$ can be obtained via sampling the corresponding features on the triplane $P_{i}$: $(c,\sigma)=f_{\text{mlp}}(P_{i},x)$. By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering [17]:
$\small\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
where $T_{r}$ is the randomly sampled viewpoints as shown in Fig. 3. $\hat{I_{r}}$ is the rendered image on view $T_{r}$, and $\hat{D_{r}}$ is the rendered depth map. Then we can use the corresponding condition extractor $f_{\text{cond}}$ introduced in Sec. 3.2 and Fig. 3 (b) to extract the condition map on the sampled view $T_{r}$:
$\small\hat{C_{r}}=f_{\text{cond}}(\hat{I_{r}}),$ (3)
where $\hat{C_{r}}$ is the extracted condition map. Note that the condition extractor $f_{\text{cond}}$ is differentiable.
Figure 4: Visualization results of our CyC3D. The input conditions and the rendered images at different views are visualized. Our CyC3D can generate high quality results with consistent conditional controls.
In a second time, we can feed the obtained condition map $\hat{C_{r}}$ on novel view $T_{r}$ back to the generation model $f_{\text{gen}}$ again:
$\small P_{r}=f_{\text{gen}}(\hat{C_{r}},C_{t},T_{r}),$ (4)
where the predicted triplane $P_{r}$ can further be used to render on the input reference view $T_{i}$:
$\small\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
where $\overline{I_{i}}$ is the rendered image on the input reference view, and $\overline{D_{i}}$ is the rendered depth map. Then we can extract the conditon map from the rendered image $\overline{I_{i}}$:
$\small\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
where $\overline{C_{i}}$ is the generated condition map back on the input reference view.
As shown in Fig. 2 and 3, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls. Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedbacks as follows:
2D Conditional Control Feedback: If we only consider 2D conditional controls extracted from the rendered images (e.g. Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal. The generated condition map $\overline{C_{i}}$ via Eq. 6 should be consistent with the original condition map $C_{v}$ on the input reference view.
$\small L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
If the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig. 3 (c)), saving the computation cost:
$\small\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
$\small\hat{C_{i}}=f_{\text{cond}}(\hat{I_{i}}),$ (9)
where $\hat{C_{i}}$ is the extracted condition map on the input reference view. In this way, we can modify Eq. 7 as:
$\small L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
Figure 5: Qualitative comparison among our CyC3D and the other state-of-the-art controllable 3D generation methods.
3D Conditional Control Feedback: If we consider 3D conditional controls like depth or normal conditions, the 3D shape of the generated objects should be regularized as well for a fine-grained control.
(1) If the input condition is depth map ($C_{v}=C_{d}$), the explicit consistency on the rendered depth map and the input conditional depth map can be calculated as follows:
$\small L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
where $f_{\text{norm}}$ is the depth normalizing function. Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following [22].
In analogy with Eq. 10, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can modify Eq. 11 as follows:
$\small L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
where $\hat{D_{i}}$ can be obtained from Eq. 8.
(2) If the input condition is normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
where $f_{\text{d2n}}$ is a differentiable module [13] that can converts the depth map to a unified normal map.
In analogy with Eq. 10 and 12, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can thus modify Eq. 13 as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)
where $\hat{D_{i}}$ can be obtained from Eq. 8.
Figure 6: Qualitative comparison of controllability among our CyC3D and the existing state-of-the-art methods in controllable 3D generation.
3.4 View Cycle Consistency Regularization
Since the condition cycle consistency regularization (Sec. 3.3) is a strong constraint that enforces the reconstructed and the input conditions to be as similar as possible, only using this loss might lead to overfitting towards the input viewpoint. If the generated 3D content from the initial view fails, the remaining processes might be misled to a wrong cross-view mapping. Only using the condition cycle consistency on the input view can not handle this issue, and the semantic multi-view consistency might be ignored during optimization. Consequently, we propose View Cycle Consistency Regularization to ensure convergence.
Rendering Regularization Loss: To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views are used to supervise the generation model $f_{\text{gen}}$:
$\small L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
where $T_{l}$ means the pre-defined ground truth viewpoints in the multi-view dataset, and $I_{l}^{\text{gt}}$ is the corresponding ground truth image. Note that this constraint can only be used in the first-time feedforward computation of the generation model. Because the randomness of the generation model might be depressed if enforcing the first-time and second-time generated 3D contents to be the same.
Semantic Consistency Loss: We utilize the text prompt $C_{t}$ to regularize the semantic consistency of different stages.
$\small L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
$(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\overline{I}_{i}))).$
View Consistency Regularization: The overall view cycle consistency loss can be computed as follows:
$\small L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
where $\alpha$ is the set to 5.0 following [31] to balance the scale of different terms as default.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
GaussianDreamer [34] 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
DreamGaussian [26] 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
VolumeDiffusion [28] 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
3DTopia [11] 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
MVControl [15] 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
ControLRM-T [31] 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
ControLRM-D [31] 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
CyC3D-T (Ours) 15.17 0.885 0.0342 18.31 0.889 0.0159 0.0051 0.0037 0.0065 0.0027 0.0062
CyC3D-D (Ours) 15.09 0.885 0.0347 18.38 0.890 0.0156 0.0051 0.0041 0.0060 0.0028 0.0060
Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
GaussianDreamer [34] 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
DreamGaussian [26] 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
VolumeDiffusion [28] 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
3DTopia [11] 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
MVControl [15] 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
ControLRM-T [31] 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
ControLRM-D [31] 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
CyC3D-T (Ours) 12.98 0.756 0.0540 16.55 0.847 0.0225 0.0080 0.0078 0.0102 0.0029 0.0102
CyC3D-D (Ours) 12.99 0.755 0.0539 16.45 0.845 0.0231 0.0075 0.0087 0.0091 0.0030 0.0109
Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
3.5 Overall Loss
Finally, the overall loss is the combination of the aforementioned losses:
$L_{\text{total}}=\left\{\begin{aligned} &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-d}},\;\;\text{if}\;\;C_{v}=C_{d}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-n}},\;\;\text{if}\;\;C_{v}=C_{n}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}},\quad\text{otherwise}\end{aligned}\right.$ (18)
where $\lambda$ is set to 1.0 in default, and $\beta$ is set to 0.1 for regularization.
4 Experiment
4.1 Implementation Details
Datasets. Our training dataset is the training split of G-Objaverse (GOBJ) dataset [21], a sub-set of Objaverse [8]. Following the selection principle of [15], we utilize the test samples collected by [31] gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) [9], and 80 samples from Amazon Berkeley Objects dataset (ABO) [7].
[…]
FID CLIP-I CLIP-T FID CLIP-I CLIP-T
GSGEN [6] 344.61 0.740 0.289 366.47 0.669 0.259
GaussianDreamer [34] 278.70 0.810 0.300 225.38 0.787 0.277
DreamGaussian [26] 359.65 0.760 0.279 392.95 0.723 0.247
VolumeDiffusion [28] 299.61 0.715 0.259 350.46 0.679 0.242
3DTopia [11] 331.39 0.727 0.283 231.55 0.751 0.272
MVControl [15] 278.08 0.816 0.288 217.97 0.802 0.291
ControLRM-T [31] 260.75 0.846 0.289 202.14 0.827 0.282
ControLRM-D [31] 171.13 0.838 0.302 181.84 0.836 0.292
CyC3D-T (Ours) 255.30 0.840 0.295 196.55 0.836 0.291
CyC3D-D (Ours) 168.71 0.846 0.303 183.84 0.842 0.293
Table 3: Generation quality (FID) and CLIP score (CLIP-I, CLIP-T) comparison with state-of-the-art controllable 3D generation methods GSO, and ABO benchmarks. $\uparrow$ means higher result is better, while $\downarrow$ means lower is better. The best results are emphasized in bold. We provide the average results of 4 different conditions (edge, sketch, depth, and normal) in the table.
Training Details. In analogy with [31], we initialize our network with the weights from pretrained OpenLRM-base [12]. For CyC3D-T, the conditional backbone is a transformer-based network; For CyC3D-D, the conditional backbone is a diffusion-based network. The training might cost 2 days for CyC3D-T and 3 days for CyC3D-D on 32 Nvidia V100-32G GPUs. More details are provided in the appendix.
Evaluation and Metrics. For the task of controllable 3D generation task, we consider 4 different conditions: edge, sketch, depth, and normal. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following [31]. For edge/sketch condition, we utilize PSNR, SSIM, and MSE as evaluation metrics; For depth/normal condition, we utilize MSE to measure the similarity. In analogy with [14], the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation. If Midas [22] is used, the metric is noted as M-MSE. If ZoeDepth [3] is used, the metric is noted as Z-MSE. Following [31], the depth consistency in 3D space compares the difference between rendered depth map and the input condition depth map, noted as R-MSE. In analogy with [14], the metric of NB-MSE uses the pre-trained model (Normal-BAE [1]) in surface normal estimation to assess the normal consistency. Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following [31], noted as DN-CON. Please refer to more detailed introduction of these evaluation metrics in the appendix.
4.2 Experimental Results
Methods Control Quality
Edge Sketch Depth Normal FID CLIP-I CLIP-T
ControlNet+OpenLRM-S [12] 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
ControlNet+OpenLRM-B [12] 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
ControlNet+OpenLRM-L [12] 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
ControlNet+TGS [36] 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
ControlNet+TripoSR [29] 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
MVControlNet+LGM [15] 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
CyC3D-T (Ours) 0.0342 0.0159 0.0065 0.0062 255.30 0.840 0.295
CyC3D-D (Ours) 0.0347 0.0156 0.0060 0.0060 168.71 0.846 0.303
Table 4: Comparison with ControlNet-based baselines (ControlNet + Image-to-3D methods) on GSO benchmark.
Qualitative Results. We provide qualitative results of our CyC3D in Fig. 4. The backbone is inherited from ControLRM [31], and the diffusion model is replaced with a pre-trained ControlNet. As the figure shows, our method not only performs well in generating hight quality 3D contents but also in preserving details in conditional controls. Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods including ControLRM [31] and MVControl [15] in Fig. 5.
Controllability Evaluation. The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab. 1/2. Following [31], we report the controllability metrics discussed in Sec. 4.1 on all conditional controls. From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines. For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE). Furthermore, we also provide qualitative comparison of controllability in Fig. 6.
Generation Evaluation. To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) [10] and CLIP-score [18] following [15] in Tab. 3. We report CLIP-I metric to quantify image similarity between reference and rendered images, and CLIP-T metric to evaluate text-image alignment using textual annotations. In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods. Furthermore, from the qualitative results in Fig. 5, out CyC3D can generate better results with higher quality and consistency than existing methods.
Comparison with ControlNet-based Baselines. For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models. The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal. The quantitative evaluation results on the GSO benchmark are presented in Tab. 4, incorporating state-of-the-art image-to-3D systems including OpenLRM [12], TGS [36], and TripoSR [29] as comparative baselines.
Loss Edge Sketch
PSNR SSIM MSE PSNR SSIM MSE
w/o $L_{\text{cond}}$ 12.87 0.841 0.0570 17.52 0.888 0.0187
w $L_{\text{cond}}$ 15.09 0.885 0.0347 18.38 0.890 0.0156
Table 5: Ablation experiments of 2D condition cycle consistency.
Loss Depth Loss Normal
M-MSE Z-MSE R-MSE NB-MSE DN-CON
w/o $L_{\text{cond}}$ 0.0383 0.0546 0.0650 w/o $L_{\text{cond}}$ 0.0034 0.0205
w $L_{\text{cond}}$ 0.0192 0.0241 0.0321 w $L_{\text{cond}}$ 0.0023 0.0187
w $L_{\text{cond}}+L_{\text{cond-d}}$ 0.0051 0.0041 0.0060 w $L_{\text{cond}}+L_{\text{cond-n}}$ 0.0011 0.0056
Table 6: Ablation experiments of 3D condition cycle consistency.
Figure 7: Ablation experiments of view cycle consistency.
4.3 Ablation Study
Effect of Condition Cycle Consistency. To evaluate the effectiveness of 2D conditional control feedback (e.g. Eq. 10 for edge/sketch), we provide ablation experiments in Tab. 5. Furthermore, we also provide ablation results of 3D conditional control feedback (Eq. 12 for depth, Eq. 14 for normal) in Tab. 6. As the tables show, each term of condition cycle consistency can effectively imporve the controllability upon different conditional controls.
Effect of View Cycle Consistency. To evaluate the effectiveness of view cycle consistency feedback (e.g. Eq. 17), we provide qualitative ablation results in Fig. 7. As the figure shows, without view cycle consistency regularization, the generated 3D contents might be overfitting to the input viewpoint. Though the rendered results in the input view is reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.
5 Conclusion
In this paper, we propose CyC3D, a novel framework for improving the controllability during 3D generation guided by different conditional controls (edge/sketch/depth/normal). The core idea is the cycle-view consistency which ensures accurate alignment between input conditions and generated models by rendering the generated 3D content from various viewpoints and re-extracting the conditions from these rendered views to feed back into the generation process. The quantitative and qualitative evaluation results on GSO and ABO benchmarks show the state-of-the-art controllability of CyC3D.

### v2

@@
 Despite the remarkable progress of 3D generation, achieving controllability, i.e., ensuring consistency between generated 3D content and input conditions like edge and depth, remains a significant challenge.
-Existing methods often struggle to maintain accurate alignment, leading to noticeable discrepancies.
-To address this issue, we propose CyC3D, a new framework that enhances controllable 3D generation by explicitly encouraging cyclic consistency between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls.
+Existing approaches often struggle to maintain accurate alignment, leading to noticeable discrepancies.
+To address this issue, we propose CyC3D, a new framework designed to enhance controllable 3D generation by explicitly encouraging cyclic consistency during training between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls.
 Specifically, we employ an efficient feed-forward backbone that can generate a 3D object from an input condition and a text prompt.
@@
 Condition consistency aligns the final extracted signal with the original input control, preserving structural or geometric details throughout the process.
-Extensive experiments on popular benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
+Extensive experiments on zero-shot GSO/ABO benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
 1 Introduction
-Figure 1: Visualization of input and reconstructed conditions.
-We feed the input condition to the generation model first and use the generated 3D contents to render the image at the input viewpoint to extract the reconstructed conditions.
-The difference between the input and the reconstructed conditions can reveal whether the fine-grained controls are reserved in the results.
 Figure 1: Brief overview of our proposed CyC3D.
-We first feed the condition on the original reference view to the generation model and extract the conditions from the rendered results on randomly sampled novel views.
-The extracted conditions are then fed to the generation model again and rendered back to the original reference view.
-The extracted condition of the rendered results should be consistent with the input condition on the reference view.
-The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR [2], digital game design [32], and filmmaking [30].
+The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR (Behravan 2025), digital game design (Xu et al.
+2024b), and filmmaking (Wang et al.
+2025).
 The mainstream approaches can be broadly categorized into two groups: optimization-based methods and feed-forward methods.
-Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) [20] to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models [24].
-In contrast, feed-forward 3D generation approaches [19, 28, 11, 12, 36, 29, 5] generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
+Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) (Poole et al.
+2022) to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models (Rombach et al.
+2022).
+In contrast, feed-forward 3D generation approaches (Nichol et al.
+2022; Tang et al.
+2023b; Hong et al.
+2024, 2023; Zou et al.
+2024; Tochilkin et al.
+2024; Chen et al.
+2025) generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
 While significant progress has been made in 3D generation research, a critical yet under-explored challenge remains—the controllability of 3D generation.
@@
 By incorporating these structured constraints, controllable 3D generation enhances the flexibility and reliability of the synthesis process, enabling more precise manipulation of the generated content.
-To advance research in controllable 3D generation, MVControl [15] extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model [25].
-It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model [27] to generate the final 3D representation.
-In contrast, ControLRM [31] aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) with a conditional encoder, enabling more precise conditioning.
-While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., achieving precise and fine-grained conditional control.
+To advance research in controllable 3D generation, MVControl (Li et al.
+2024) extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model (Shi et al.
+2023).
+It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model (Tang et al.
+2025a) to generate the final 3D representation.
+In contrast, ControLRM (Xu et al.
+2024a) aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) (Hong et al.
+2023) with a conditional encoder, enabling more precise conditioning.
+The inference pipeline of these controllable 3D generation works is summarized in Fig.
+1 (a).
+While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., , achieving precise and fine-grained conditional control.
 As illustrated in Fig.
-1, the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
+1 (b), the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
 To address this challenge, we propose CyC3D, a novel and general training framework that enhances the controllability of 3D generation.
 Our approach extends a feed-forward 3D generation network into a cyclic paradigm, enabling more precise and consistent adherence to input control signals, as illustrated in Fig.
-2.
+1 (c).
 Specifically, given an input control condition on a reference view, the generated 3D content is first rendered from randomly sampled new viewpoints.
@@
 To further refine controllability, we introduce two complementary consistency constraints: View Cycle Consistency, which ensures that the two generated contents have similar semantics in embedded space since they are generated with the same prompt and highly coupled conditions, and Condition Cycle Consistency, which enforces strict alignment between the extracted control signals from the final rendered images and the original input condition.
-We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets [31] for zero-shot evaluation.
+We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets (Xu et al.
+2024a) for zero-shot evaluation.
 Experimental results demonstrate that our framework significantly enhances controllability across all benchmarks while preserving competitive generation quality compared to existing methods.
-We highlight the contributions of this work as follows:
-  • New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency (Fig.
-2).
-  • Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency.
+Figure 2: Training framework of our proposed CyC3D. It contains two kinds of regularization: condition cycle consistency regularization (Sec. 3.3) and view cycle consistency regularization (Sec. 3.4).
+We highlight the contributions of this work as follows: (1) New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency.
+(2) Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency.
 The former regularizes the consistency between input and reconstructed conditions, the latter enhances the semantic consistency of different stages.
-  • Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
+(3) Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
 2 Related Work
@@
 3D generation can be broadly categorized into optimization-based and feed-forward approaches: (1) Optimization-based methods offer a viable alternative to traditional techniques by circumventing the need for extensive text-to-3D datasets.
-DreamFusion [20] exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets.
+DreamFusion (Poole et al.
+2022) exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets.
 Despite their effectiveness, these methods often encounter the Janus problem, which arises from inconsistencies between 2D supervision and 3D geometry.
 (2) Feed-forward models have emerged as a prominent approach for efficient 3D generation, leveraging extensive 3D datasets.
-Techniques such as LRM [12] and TripoSR [29] utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities.
-Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer [16] and MVDream [25].
+Techniques such as LRM (Hong et al.
+2023) and TripoSR (Tochilkin et al.
+2024) utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities.
+Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer (Liu et al.
+2023) and MVDream (Shi et al.
+2023).
+Though Cycle3D (Tang et al.
+2025b) builds a cycle-consistent pipeline of diffusion and reconstruction models, the whole pipeline is built in the inference model.
+Whereas our cycle-consistency is used for training a real-time feed-forward generation model.
 Controllable 3D Generation.
@@
 Techniques developed for controllable 2D generation have provided valuable insights and methods that are applicable to the more complex domain of controllable 3D generation.
-Inspired by ControlNet [35], MVControl [15] integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation.
+Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), MVControl (Li et al.
+2024) integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation.
 This method combines coarse Gaussian generation with SDS-based refinement but faces challenges related to model generalization and procedural complexity.
-ControLRM [31] addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation.
-Building on the advances of ControlNet++ [14] and ControLRM [35], we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
-3 CyC3D
-Figure 3: Overall framework of our proposed CyC3D.
-(a) summarize the procedure of our CyC3D.
-(b) shows the condition extractor utilized for different conditional controls (Sec.
-3.2).
-(c) illustrates the condition cycle consistency regularization (Sec.
-3.3).
-(d) introduces the proposed view cycle consistency regularization (Sec.
-3.4).
-As shown in Fig.
-3, we present the overall framework of our proposed CyC3D.
-In Sec.
-3.1, we first introduce the feed-forward backbone network for controllable 3D generation in our CyC3D.
-Then we introduce the condition extractor (Fig.
-3(b)) for different conditional controls (e.g.
-edge, sketch, depth, normal) in Sec.
-3.2.
-In Sec.
-3.3, we introduce the design of condition cycle consistency regularization (Fig.
-3(c)) to enhance the fined-grained controls.
-In Sec.
-3.4, we further introduce the proposed view cycle consistency regularization (Fig.
-3(d)) to stabilize the training and avoid overfitting on the input condition.
-Finally, we introduce the overall loss in Sec.
-3.5.
+ControLRM (Xu et al.
+2024a) addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation.
+Building on the advances of ControlNet++ (Li et al.
+2025) and ControLRM (Zhang, Rao, and Agrawala 2023), we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
+3 Method
 3.1 Generation Backbone
 For controllable 3D generation, our CyC3D can be combined with arbitrary feed-forward backbones (e.g.
-ControLRM [31], MVControl [15]).
-In default, we select ControLRM [31] as our backbone model.
+ControLRM (Xu et al.
+2024a), MVControl (Li et al.
+2024)).
+In default, we select ControLRM (Xu et al.
+2024a) as our backbone model.
 ControLRM contains a 2D condition extractor and a 3D triplane decoder.
-(1) The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs.
+The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs.
 This generator can utilize either a transformer or diffusion backbone to create initial 2D latent representations, which are subsequently transformed into 3D models.
-(2) The 3D triplane decoder is inherited from a pre-trained large reconstruction model [12].
+The 3D triplane decoder is inherited from a pre-trained large reconstruction model (Hong et al.
+2023).
 The triplane transformer decoder employs a unique representation of 3D data by decomposing it into three orthogonal planes (X, Y, and Z).
 Each plane is processed using transformer layers, which utilize self-attention mechanisms to capture spatial relationships and dependencies within the data.
-3.2 Condition Extractor
-To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor in Fig.
-2 should be differentiable.
+Figure 3: Visualization results of our CyC3D.
+The input conditions and the rendered images at different views are visualized.
+Our CyC3D can generate high-quality results with consistent conditional controls.
+3.2 Condition Extraction
+To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor should be fully differentiable.
 The design of the condition extractor for each condition control is shown as follows:
-Canny Condition. As shown in the figure, we extract edge maps using the Canny algorithm [4]. Following [14], a differentiable Canny extractor can implemented based on Kornia [23] to extract edge maps.
+Canny Condition.
+We extract edge maps using the Canny algorithm (Canny 1986).
+Following (Li et al.
+2025), a differentiable Canny extractor can be implemented based on Kornia (Riba et al.
+2020) to extract edge maps.
 Sketch Condition.
-Following ControlNet [35], we utilize a pre-trained CNN network to extract sketch conditions from images.
-Since the CNN network is differentiable, the gradients can be passed backward through the network and back-propagated to the remaining modules.
+Following ControlNet (Zhang, Rao, and Agrawala 2023), we utilize a pre-trained CNN network to extract sketch conditions from images.
+Since the CNN network is differentiable, the gradients can be passed backward through the network.
 Depth Condition.
-In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything [33]) to extract the depth condition from rendered images.
-The whole process is differentiable in the cycle consistency pipeline.
+In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything (Yang et al.
+2024)) to extract the depth condition from rendered images.
+As an alternative, we can also normalize the rendered depth map via volume rendering to extract the geometric prior from the 3D content.
 Normal Condition.
-In practice, we find that the existing normal estimation models [1] utilized in ControlNet [35] tends to predict over-smoothed normal maps.
+In practice, we find that the existing normal estimation models (Bae, Budvytis, and Cipolla 2021) utilized in ControlNet (Zhang, Rao, and Agrawala 2023) tend to predict over-smoothed normal maps.
 It may result in a large gap between the exact 3D shape of the rendered multi-view images and the estimated normal maps.
-Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to normal map via a differentiable module [13].
-The great zero-shot generalization ability of large depth-estimation models like Depth-Anything [33] can effectively capture the 3D prior from the image and thus generate accurate normal map.
+Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to a normal map via a differentiable module (Huang et al.
+2021).
+As an alternative, we also extract the normal from the rendered depth maps via volume rendering.
 3.3 Condition Cycle Consistency Regularization
@@
 Given the text prompt $C_{t}$ and the input visual condition as $C_{v}$, the generation model outputs the triplane representation $P\in\mathbb{R}^{3\times C_{p}\times H_{p}\times W_{p}}$ corresponding to the input conditions.
-$\small P_{i}=f_{\text{gen}}(C_{v},C_{t},T_{i}),$ (1)
+$P_{i}=f_{\text{gen}}(C_{v},C_{t},T_{i}),$ (1)
 where $T_{i}$ is the camera pose of the input condition, which is usually assumed to be an identity matrix after normalizing all the camera extrinsics to a frontal view.
@@
 Given a 3D point $x\in\mathbb{R}^{3}$, the color and density filed $(c,\sigma)$ can be obtained via sampling the corresponding features on the triplane $P_{i}$: $(c,\sigma)=f_{\text{mlp}}(P_{i},x)$.
-By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering [17]:
-$\small\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
+By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering (Mildenhall et al.
+2021):
+$\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
 where $T_{r}$ is the randomly sampled viewpoints as shown in Fig.
-3.
+2.
 $\hat{I_{r}}$ is the rendered image on view $T_{r}$, and $\hat{D_{r}}$ is the rendered depth map.
@@
 2 (b) to extract the condition map on the sampled view $T_{r}$:
-$\small\hat{C_{r}}=f_{\text{cond}}(\hat{I_{r}}),$ (3)
+[…]
 where $\hat{C_{r}}$ is the extracted condition map. Note that the condition extractor $f_{\text{cond}}$ is differentiable.
-Figure 4: Visualization results of our CyC3D.
-The input conditions and the rendered images at different views are visualized.
-Our CyC3D can generate high quality results with consistent conditional controls.
 In a second time, we can feed the obtained condition map $\hat{C_{r}}$ on novel view $T_{r}$ back to the generation model $f_{\text{gen}}$ again:
-$\small P_{r}=f_{\text{gen}}(\hat{C_{r}},C_{t},T_{r}),$ (4)
+$P_{r}=f_{\text{gen}}(\hat{C_{r}},C_{t},T_{r}),$ (4)
 where the predicted triplane $P_{r}$ can further be used to render on the input reference view $T_{i}$:
-$\small\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
+$\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
 where $\overline{I_{i}}$ is the rendered image on the input reference view, and $\overline{D_{i}}$ is the rendered depth map.
-Then we can extract the conditon map from the rendered image $\overline{I_{i}}$:
-$\small\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
+Then we can extract the condition map from the rendered image $\overline{I_{i}}$:
+$\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
 where $\overline{C_{i}}$ is the generated condition map back on the input reference view.
 As shown in Fig.
-2 and 3, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls.
-Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedbacks as follows:
-2D Conditional Control Feedback: If we only consider 2D conditional controls extracted from the rendered images (e.g.
-Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal.
+2, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls.
+Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedback as follows:
+2D Conditional Control Feedback.
+If we only consider 2D conditional controls extracted from the rendered images (e.g., Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal.
 The generated condition map $\overline{C_{i}}$ via Eq.
 6 should be consistent with the original condition map $C_{v}$ on the input reference view.
-$\small L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
-If the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig.
+$L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
+If the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig.
 2 (c)), saving the computation cost:
-$\small\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
-$\small\hat{C_{i}}=f_{\text{cond}}(\hat{I_{i}}),$ (9)
+$\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
+[…]
 where $\hat{C_{i}}$ is the extracted condition map on the input reference view. In this way, we can modify Eq. 7 as:
-$\small L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
+$L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
 Figure 4: Qualitative comparison among our CyC3D and the other state-of-the-art controllable 3D generation methods.
-3D Conditional Control Feedback: If we consider 3D conditional controls like depth or normal conditions, the 3D shape of the generated objects should be regularized as well for a fine-grained control.
+In the 1st row, the baseline methods fail to generate consistent content with the text prompts.
+In the 2nd row, ControLRM generates a fainter color compared with our CyC3D.
+3D Conditional Control Feedback. If we consider 3D conditional controls, such as depth or normal conditions, the 3D shape of the generated objects should also be regularized for fine-grained control.
 (1) If the input condition is depth map ($C_{v}=C_{d}$), the explicit consistency on the rendered depth map and the input conditional depth map can be calculated as follows:
-$\small L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
+$L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
 where $f_{\text{norm}}$ is the depth normalizing function.
-Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following [22].
+Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following (Ranftl et al.
+2020).
 In analogy with Eq.
-10, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation.
+10, when the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation.
 We can modify Eq.
 11 as follows:
-$\small L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
+$L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
 where $\hat{D_{i}}$ can be obtained from Eq. 8.
-(2) If the input condition is normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
-$\small L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
-where $f_{\text{d2n}}$ is a differentiable module [13] that can converts the depth map to a unified normal map.
+(2) If the input condition is a normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
+$L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
+where $f_{\text{d2n}}$ is a differentiable module (Huang et al. 2021) that can convert the depth map to a unified normal map.
 In analogy with Eq.
-10 and 12, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation.
+10 and 12, when the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation.
 We can thus modify Eq.
 13 as follows:
-$\small L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)
+$L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)
 where $\hat{D_{i}}$ can be obtained from Eq. 8.
+Methods Edge Sketch Depth Normal
+PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
+GSGEN (Chen et al. 2024) 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
+GaussianDreamer (Yi et al. 2024) 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
+DreamGaussian (Tang et al. 2023a) 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
+VolumeDiffusion (Tang et al. 2023b) 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
+3DTopia (Hong et al. 2024) 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
+MVControl (Li et al. 2024) 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
+ControLRM-T (Xu et al. 2024a) 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
+ControLRM-D (Xu et al. 2024a) 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
+CyC3D-T (Ours) 15.17 0.885 0.0342 18.31 0.889 0.0159 0.0051 0.0037 0.0065 0.0027 0.0062
+CyC3D-D (Ours) 15.09 0.885 0.0347 18.38 0.890 0.0156 0.0051 0.0041 0.0060 0.0028 0.0060
+Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark.
+$\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better.
+The best results are highlighted in bold.
+We provide the results of four different conditions: edge, sketch, depth, and normal.
+Methods Edge Sketch Depth Normal
+PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
+GSGEN (Chen et al. 2024) 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
+GaussianDreamer (Yi et al. 2024) 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
+DreamGaussian (Tang et al. 2023a) 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
+VolumeDiffusion (Tang et al. 2023b) 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
+3DTopia (Hong et al. 2024) 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
+MVControl (Li et al. 2024) 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
+ControLRM-T (Xu et al. 2024a) 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
+ControLRM-D (Xu et al. 2024a) 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
+CyC3D-T (Ours) 12.98 0.756 0.0540 16.55 0.847 0.0225 0.0080 0.0078 0.0102 0.0029 0.0102
+CyC3D-D (Ours) 12.99 0.755 0.0539 16.45 0.845 0.0231 0.0075 0.0087 0.0091 0.0030 0.0109
+Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark.
+$\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better.
+The best results are highlighted in bold.
+We provide the results of four different conditions: edge, sketch, depth, and normal.
 Figure 5: Qualitative comparison of controllability among our CyC3D and the existing state-of-the-art methods in controllable 3D generation.
@@
 Consequently, we propose View Cycle Consistency Regularization to ensure convergence.
-Rendering Regularization Loss: To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views are used to supervise the generation model $f_{\text{gen}}$:
-$\small L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
+Rendering Regularization Loss.
+To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views is used to supervise the generation model $f_{\text{gen}}$:
+$L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
 where $T_{l}$ means the pre-defined ground truth viewpoints in the multi-view dataset, and $I_{l}^{\text{gt}}$ is the corresponding ground truth image.
@@
 Semantic Consistency Loss: We utilize the text prompt $C_{t}$ to regularize the semantic consistency of different stages.
-$\small L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
+$L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
 $(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\overline{I}_{i}))).$
 View Consistency Regularization: The overall view cycle consistency loss can be computed as follows:
-$\small L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
-where $\alpha$ is the set to 5.0 following [31] to balance the scale of different terms as default.
-Methods Edge Sketch Depth Normal
-PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
-GSGEN [6] 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
-GaussianDreamer [34] 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
-DreamGaussian [26] 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
-VolumeDiffusion [28] 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
-3DTopia [11] 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
-MVControl [15] 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
-ControLRM-T [31] 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
-ControLRM-D [31] 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
-CyC3D-T (Ours) 15.17 0.885 0.0342 18.31 0.889 0.0159 0.0051 0.0037 0.0065 0.0027 0.0062
-CyC3D-D (Ours) 15.09 0.885 0.0347 18.38 0.890 0.0156 0.0051 0.0041 0.0060 0.0028 0.0060
-Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark.
-$\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better.
-The best results are highlighted in bold.
-We provide the results of four different conditions: edge, sketch, depth, and normal.
-Our CyC3D-T/-D achieves significant controllability improvements in the table.
-Methods Edge Sketch Depth Normal
-PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
-GSGEN [6] 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
-GaussianDreamer [34] 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
-DreamGaussian [26] 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
-VolumeDiffusion [28] 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
-3DTopia [11] 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
-MVControl [15] 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
-ControLRM-T [31] 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
-ControLRM-D [31] 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
-CyC3D-T (Ours) 12.98 0.756 0.0540 16.55 0.847 0.0225 0.0080 0.0078 0.0102 0.0029 0.0102
-CyC3D-D (Ours) 12.99 0.755 0.0539 16.45 0.845 0.0231 0.0075 0.0087 0.0091 0.0030 0.0109
-Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark.
-$\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better.
-The best results are highlighted in bold.
-We provide the results of four different conditions: edge, sketch, depth, and normal.
-Our CyC3D-T/-D achieves significant controllability improvements in the table.
+$L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
+where $\alpha$ is the set to 5.0 following (Xu et al. 2024a) to balance the scale of different terms as default.
 3.5 Overall Loss
 Finally, the overall loss is the combination of the aforementioned losses:
-$L_{\text{total}}=\left\{\begin{aligned} &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-d}},\;\;\text{if}\;\;C_{v}=C_{d}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-n}},\;\;\text{if}\;\;C_{v}=C_{n}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}},\quad\text{otherwise}\end{aligned}\right.$ (18)
+$L_{\text{total}}=\left\{\begin{aligned} &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-d}},\;\;\text{if}\;\;C_{v}=C_{d}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}}+\beta\cdot L_{\text{cond-n}},\;\;\text{if}\;\;C_{v}=C_{n}\\ &L_{\text{view}}+\lambda\cdot L_{\text{cond}},\quad\text{otherwise}\end{aligned}\right$.
+(18)
 where $\lambda$ is set to 1.0 in default, and $\beta$ is set to 0.1 for regularization.
@@
 Datasets.
-Our training dataset is the training split of G-Objaverse (GOBJ) dataset [21], a sub-set of Objaverse [8].
-Following the selection principle of [15], we utilize the test samples collected by [31] gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) [9], and 80 samples from Amazon Berkeley Objects dataset (ABO) [7].
-[…]
+Our training dataset is the training split of the G-Objaverse (GOBJ) dataset (Qiu et al.
+2024), a sub-set of Objaverse (Deitke et al.
+2023).
+Following the selection principle of (Li et al.
+2024), we utilize the test samples collected by (Xu et al.
+2024a) gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) (Downs et al.
+2022), and 80 samples from Amazon Berkeley Objects dataset (ABO) (Collins et al.
+2022).
+Methods $\rightarrowGSO GOBJ\rightarrowABO$
 FID CLIP-I CLIP-T FID CLIP-I CLIP-T
-GSGEN [6] 344.61 0.740 0.289 366.47 0.669 0.259
-GaussianDreamer [34] 278.70 0.810 0.300 225.38 0.787 0.277
-DreamGaussian [26] 359.65 0.760 0.279 392.95 0.723 0.247
-VolumeDiffusion [28] 299.61 0.715 0.259 350.46 0.679 0.242
-3DTopia [11] 331.39 0.727 0.283 231.55 0.751 0.272
-MVControl [15] 278.08 0.816 0.288 217.97 0.802 0.291
-ControLRM-T [31] 260.75 0.846 0.289 202.14 0.827 0.282
-ControLRM-D [31] 171.13 0.838 0.302 181.84 0.836 0.292
+GSGEN (Chen et al. 2024) 344.61 0.740 0.289 366.47 0.669 0.259
+GaussianDreamer (Yi et al. 2024) 278.70 0.810 0.300 225.38 0.787 0.277
+DreamGaussian (Tang et al. 2023a) 359.65 0.760 0.279 392.95 0.723 0.247
+VolumeDiffusion (Tang et al. 2023b) 299.61 0.715 0.259 350.46 0.679 0.242
+3DTopia (Hong et al. 2024) 331.39 0.727 0.283 231.55 0.751 0.272
+MVControl (Li et al. 2024) 278.08 0.816 0.288 217.97 0.802 0.291
+ControLRM-T (Xu et al. 2024a) 260.75 0.846 0.289 202.14 0.827 0.282
+ControLRM-D (Xu et al. 2024a) 171.13 0.838 0.302 181.84 0.836 0.292
 CyC3D-T (Ours) 255.30 0.840 0.295 196.55 0.836 0.291
-CyC3D-D (Ours) 168.71 0.846 0.303 183.84 0.842 0.293
+CyC3D-D (Ours) 168.71 0.846 0.303 180.34 0.843 0.294
 Table 3: Generation quality (FID) and CLIP score (CLIP-I, CLIP-T) comparison with state-of-the-art controllable 3D generation methods GSO, and ABO benchmarks.
@@
 Training Details.
-In analogy with [31], we initialize our network with the weights from pretrained OpenLRM-base [12].
+In analogy with (Xu et al.
+2024a), we initialize our network with the weights from pretrained OpenLRM-base (Hong et al.
+2023).
 For CyC3D-T, the conditional backbone is a transformer-based network; For CyC3D-D, the conditional backbone is a diffusion-based network.
@@
 For the task of controllable 3D generation task, we consider 4 different conditions: edge, sketch, depth, and normal.
-For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following [31].
+For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following (Xu et al.
+2024a).
 For edge/sketch condition, we utilize PSNR, SSIM, and MSE as evaluation metrics; For depth/normal condition, we utilize MSE to measure the similarity.
-In analogy with [14], the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation.
-If Midas [22] is used, the metric is noted as M-MSE.
-If ZoeDepth [3] is used, the metric is noted as Z-MSE.
-Following [31], the depth consistency in 3D space compares the difference between rendered depth map and the input condition depth map, noted as R-MSE.
-In analogy with [14], the metric of NB-MSE uses the pre-trained model (Normal-BAE [1]) in surface normal estimation to assess the normal consistency.
-Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following [31], noted as DN-CON.
-Please refer to more detailed introduction of these evaluation metrics in the appendix.
+In analogy with (Li et al.
+2025), the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation.
+If Midas (Ranftl et al.
+2020) is used, the metric is noted as M-MSE.
+If ZoeDepth (Bhat et al.
+2023) is used, the metric is noted as Z-MSE.
+Following (Xu et al.
+2024a), the depth consistency in 3D space compares the difference between the rendered depth map and the input condition depth map, noted as R-MSE.
+In analogy with (Li et al.
+2025), the metric of NB-MSE uses the pre-trained model (Normal-BAE (Bae, Budvytis, and Cipolla 2021)) in surface normal estimation to assess the normal consistency.
+Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following (Xu et al.
+2024a), noted as DN-CON.
+Please refer to more details in the appendix.
 4.2 Experimental Results
@@
 Edge Sketch Depth Normal FID CLIP-I CLIP-T
-ControlNet+OpenLRM-S [12] 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
-ControlNet+OpenLRM-B [12] 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
-ControlNet+OpenLRM-L [12] 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
-ControlNet+TGS [36] 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
-ControlNet+TripoSR [29] 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
-MVControlNet+LGM [15] 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
+ControlNet+OpenLRM-S (Hong et al. 2023) 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
+ControlNet+OpenLRM-B (Hong et al. 2023) 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
+ControlNet+OpenLRM-L (Hong et al. 2023) 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
+ControlNet+TGS (Zou et al. 2024) 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
+ControlNet+TripoSR (Tochilkin et al. 2024) 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
+MVControlNet+LGM (Li et al. 2024) 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
 CyC3D-T (Ours) 0.0342 0.0159 0.0065 0.0062 255.30 0.840 0.295
@@
 Table 4: Comparison with ControlNet-based baselines (ControlNet + Image-to-3D methods) on GSO benchmark.
-Qualitative Results.
-We provide qualitative results of our CyC3D in Fig.
-4.
-The backbone is inherited from ControLRM [31], and the diffusion model is replaced with a pre-trained ControlNet.
-As the figure shows, our method not only performs well in generating hight quality 3D contents but also in preserving details in conditional controls.
-Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods including ControLRM [31] and MVControl [15] in Fig.
-5.
-Controllability Evaluation.
-The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab.
-1/2.
-Following [31], we report the controllability metrics discussed in Sec.
-4.1 on all conditional controls.
-From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines.
-For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE).
-Furthermore, we also provide qualitative comparison of controllability in Fig.
-6.
-Generation Evaluation.
-To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) [10] and CLIP-score [18] following [15] in Tab.
-3.
-We report CLIP-I metric to quantify image similarity between reference and rendered images, and CLIP-T metric to evaluate text-image alignment using textual annotations.
-In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods.
-Furthermore, from the qualitative results in Fig.
-5, out CyC3D can generate better results with higher quality and consistency than existing methods.
-Comparison with ControlNet-based Baselines.
-For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models.
-The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal.
-The quantitative evaluation results on the GSO benchmark are presented in Tab.
-4, incorporating state-of-the-art image-to-3D systems including OpenLRM [12], TGS [36], and TripoSR [29] as comparative baselines.
 Loss Edge Sketch
@@
 Table 6: Ablation experiments of 3D condition cycle consistency.
+Qualitative Results.
+We provide qualitative results of our CyC3D in Fig.
+10.
+The backbone is inherited from ControLRM (Xu et al.
+2024a), and the diffusion model is replaced with a pre-trained ControlNet.
+As the figure shows, our method not only performs well in generating high-quality 3D content but also in preserving details in conditional controls.
+Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods, including ControLRM (Xu et al.
+2024a) and MVControl (Li et al.
+2024) in Fig.
+4.
+Controllability Evaluation.
+The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab.
+1/2.
+Following (Xu et al.
+2024a), we report the controllability metrics discussed in Sec.
+4.1 on all conditional controls.
+From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines.
+For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE).
+Furthermore, we also provide a qualitative comparison of controllability in Fig.
+5.
+Generation Evaluation.
+To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) (Heusel et al.
+2017) and CLIP-score (Mohammad Khalid et al.
+2022) following (Li et al.
+2024) in Tab.
+3.
+Furthermore, the qualitative results are provided in Fig.
+4.
+We report the CLIP-I metric to quantify image similarity between reference and rendered images, and the CLIP-T metric to evaluate text-image alignment using textual annotations.
+In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods.
+Comparison with ControlNet-based Baselines.
+For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models.
+The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal.
+The quantitative evaluation results on the GSO benchmark are presented in Tab.
+4, incorporating state-of-the-art image-to-3D systems including OpenLRM (Hong et al.
+2023), TGS (Zou et al.
+2024), and TripoSR (Tochilkin et al.
+2024) as comparative baselines.
 Figure 6: Ablation experiments of view cycle consistency.
@@
 Effect of Condition Cycle Consistency.
-To evaluate the effectiveness of 2D conditional control feedback (e.g.
-Eq.
+To evaluate the effectiveness of 2D conditional control feedback (e.g., Eq.
 10 for edge/sketch), we provide ablation experiments in Tab.
@@
 6.
-As the tables show, each term of condition cycle consistency can effectively imporve the controllability upon different conditional controls.
+As the tables show, each term of condition cycle consistency can effectively improve the controllability upon different conditional controls.
 Effect of View Cycle Consistency.
-To evaluate the effectiveness of view cycle consistency feedback (e.g.
-Eq.
+To evaluate the effectiveness of view cycle consistency feedback (e.g., Eq.
 17), we provide qualitative ablation results in Fig.
-7.
+6.
 As the figure shows, without view cycle consistency regularization, the generated 3D contents might be overfitting to the input viewpoint.
-Though the rendered results in the input view is reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.
+Though the rendered results in the input view are reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.
 5 Conclusion
 In this paper, we propose CyC3D, a novel framework for improving the controllability during 3D generation guided by different conditional controls (edge/sketch/depth/normal).
-The core idea is the cycle-view consistency which ensures accurate alignment between input conditions and generated models by rendering the generated 3D content from various viewpoints and re-extracting the conditions from these rendered views to feed back into the generation process.
+The core idea is the cycle-view consistency, which ensures accurate alignment between input conditions and generated models by rendering the generated 3D content from various viewpoints and re-extracting the conditions from these rendered views to feed back into the generation process.
 The quantitative and qualitative evaluation results on GSO and ABO benchmarks show the state-of-the-art controllability of CyC3D.
+Appendix
+Due to the page limit, we present additional details here as a comprehensive extension to the manuscript, including:
+  1. 1.
+    Detailed introduction of utilized training dataset, and the benchmark for evaluation. (Sec. Utilized Datasets)
+  2. 2.
+    Detailed explanation of the evaluation metric on controllability. (Sec. Evaluation Metric on Controllability)
+  3. 3.
+    Detailed explanation of the evaluation metric on generation quality. (Sec. Evaluation Metric on Generation Quality)
+  4. 4.
+    Details for reproducing the baseline results used in the paper. (Sec. Baselines for Comparisons)
+  5. 5.
+    Implementation details of our proposed CyC3D. (Sec. Implementation Details)
+  6. 6.
+    Additional experimental results including quantitative and qualitative results. (Sec. Additional Experimental Results)
+  7. 7.
+    Discussion about failure cases and limitations of the proposed method. (Sec. Discussion)
+Utilized Datasets
+Training Dataset Instead of using the full version of Objaverse (Deitke et al.
+2023) as the training dataset, we utilize G-buffer Objaverse (G-Objaverse) (Qiu et al.
+2024) as the training dataset.
+The original Objaverse contains about 800K 3D objects, while the G-Objaverse contains about 280K 3D objects.
+Same as controllable 2D generation (i.e.
+ControlNet (Zhang, Rao, and Agrawala 2023)), the training does not have to be the full Objaverse dataset, our method can work on a smaller training dataset (G-Objaverse).
+In G-Objaverse, each 3D object is rendered to RGB, Depth, and Normal maps on 40 different viewpoints, including 24 views at elevation range from $5^{\circ}$ to $30^{\circ}$, rotation $\{r\times 15^{\circ}|r\in[0,23]\}$, and 12 views at elevation from $5^{\circ}$ to $5^{\circ}$, rotation $\{r\ times30^{\circ}|r\in[0,11]\}$, and 2 views for top and bottom respectively.
+The rendered RGB images on 40 different views are used in the training dataset.
+The corresponding text caption of each 3D object is also provided by Cap3D (Luo et al.
+2024).
+After filtering the unavailable samples, we can obtain a training set with 260K samples and a test set with 2800 samples.
+Evaluation Dataset of GOBJ For evaluation, 118 samples in the separated test set of G-Objaverse are selected following (Xu et al.
+2024a), which are absent from the training data.
+The samples are collected following selection principles noted in MVControl (Li et al.
+2024) and TripoSR (Tochilkin et al.
+2024).
+The text prompts are from Cap3D (Luo et al.
+2024).
+One reference view is manually selected from all 40 available views and the condition map of edge/sketch/depth/normal is extracted from the reference image.
+The reference condition map is treated as input, and the remaining images are treated as ground truth images for evaluation.
+This evaluation set is noted as GOBJ.
+Evaluation Dataset of GSO Furthermore, 80 samples of the Google Scanned Objects dataset (Downs et al.
+2022) are selected following (Xu et al.
+2024a) for zero-shot evaluation.
+This dataset features more than one thousand 3D-scanned household items serving as a valuable resource for assessing the zero-shot generalization abilities of the proposed method.
+In analogy with GOBJ, we manually select the reference view from the overall 32 available views provided in the dataset.
+The condition map of edge/sketch/depth/normal is extracted from the reference view image.
+The text prompts are obtained via captioning the reference image with BLIP2 (Li et al.
+2023).
+The condition map and the text prompt of each sample are input to the model, and the remaining views are utilized as the ground truth images for evaluation.
+This evaluation set is noted as GSO.
+Evaluation Dataset of ABO For zero-shot evaluation, we also select 80 samples of Amazon Berkeley Objects dataset (Collins et al.
+2022) following (Xu et al.
+2024a).
+The Amazon Berkeley Objects dataset is a comprehensive 3D dataset comprising product catalog images, metadata, and artist-designed 3D models featuring intricate geometries and materials based on real household objects.
+In analogy with GSO, we manually select the reference view from the overall 72 views provided in the dataset.
+The condition map of edge/sketch/depth/normal is extracted from the selected reference view image.
+The text prompts are generated via captioning the reference image with BLIP2 (Li et al.
+2023).
+The condition map and the text prompt of each sample are input to the model, and the remaining views are utilized as the ground truth images for evaluation.
+This evaluation set is noted as ABO.
+Evaluation Metric on Controllability
+In this section, we introduce the details of the evaluation metrics of controllability on different conditional controls.
+Assume that the generated 3D content is $P$, and the camera transformation matrix is $T_{r}$, the rendered image and depth via volume rendering (Mildenhall et al. 2021) can be written as follows:
+$I_{r}$ $=f_{\text{render}}^{I}(P,T_{r})$ (19)
+$D_{r}$ $=f_{\text{render}}^{D}(P,T_{r})$
+where $I_{r}$ is the rendered image at viewpoint $r$, and $D_{r}$ is the rendered depth at viewpoint $r$.
+$f_{\text{render}}^{I}$ is the volume rendering function of the RGB image based on color and density.
+$f_{\text{render}}^{D}$ is the volume rendering function of the depth map based on density only.
+To evaluate the controllability between rendered results and input condition maps, we need a corresponding condition extractor $f_{\text{cond}}$ according to each of the conditional controls:
+$f_{\text{cond}}=\left\{\begin{aligned} &f_{\text{canny}},\;\;\text{if}\;\;C_{v}=C_{e}\;(\text{edge condition})\\ &f_{\text{sketch}},\;\;\text{if}\;\;C_{v}=C_{s}\;(\text{sketch condition})\\ &f_{\text{depth}},\;\;\text{if}\;\;C_{v}=C_{d}\;(\text{depth condition})\\ &f_{\text{normal}},\;\;\text{if}\;\;C_{v}=C_{n}\;(\text{normal condition})\end{aligned}\right$.
+(20)
+where $f_{\text{canny}}$, $f_{\text{sketch}}$, $f_{\text{depth}}$, $f_{\text{normal}}$ are respectively the condition extractor of under edge/sketch/depth/normal controls.
+$C_{v}$ is the input condition.
+$C_{e}$$C_{s}$$C_{d}$$C_{n}$ respectively represents the edge/sketch/depth/normal condition map.
+A brief summarization of controllability metrics on each condition is shown as follows:
+Controllability on Edge/Sketch
+To evaluate the controllability of visual conditional controls like edge or sketch, we modify the widely used metric of PSNR, SSIM, MSE to measure the difference between the condition maps of the input sample and generated sample.
+Note that $f_{\text{cond}}=f_{\text{canny}}$ if the conditional control is edge, and $f_{\text{cond}}=f_{\text{sketch}}$ if the conditional control is sketch.
+$f_{\text{canny}}$ is the Canny descriptor (Canny 1986) which extracts edge maps from the input images.
+We share the same hyper-parameter setting of $f_{\text{canny}}$ as ControlNet (Zhang, Rao, and Agrawala 2023).
+$f_{\text{sketch}}$ is a pre-trained sketch extraction network provided by ControlNet (Zhang, Rao, and Agrawala 2023).
+$C_{\text{cond}}$ is the input conditional map at viewpoint $r$.
+$C_{\text{cond}}=C_{e}$ if the conditional control is edge, and $C_{\text{cond}}=C_{s}$ if the conditional control is sketch.
+In the following, we provide the definition of the quantitative metrics for evaluating the controllability:
+PSNR Given the camera pose $T_{r}$ at viewpoint $r$, the Peak signal-to-noise ratio (PSNR) is calculated between the condition map of the rendered image in the generated 3D content, and the one of the ground truth image $I_{r}^{\text{gt}})$ at viewpoint $r$:
+$\text{PSNR}_{\text{cond}}(P,T_{r},I_{r}^{\text{gt}})=\text{PSNR}(f_{\text{cond}}(f_{\text{render}}^{I}(P,T_{r})),C_{\text{cond}})$ (21)
+SSIM Given the camera pose $T_{r}$ at viewpoint $r$, the Structure Similarity Index Measure (SSIM) is calculated between the condition map of the rendered image in the generated 3D content, and the one of the ground truth image $I_{r}^{\text{gt}})$ at viewpoint $r$:
+$\text{SSIM}_{\text{cond}}(P,T_{r},I_{r}^{\text{gt}})=\text{SSIM}(f_{\text{cond}}(f_{\text{render}}^{I}(P,T_{r})),C_{\text{cond}})$ (22)
+MSE Given the camera pose $T_{r}$ at viewpoint $r$, the Mean Squared Error (MSE) difference is calculated between the condition map of the rendered image in the generated 3D content, and the one of the ground truth image $I_{r}^{\text{gt}})$ at viewpoint $r$:
+$\text{MSE}_{\text{cond}}(P,T_{r},I_{r}^{\text{gt}})=\text{MSE}(f_{\text{cond}}(f_{\text{render}}^{I}(P,T_{r})),C_{\text{cond}})$ (23)
+Controllability on Depth
+To evaluate the controllability of depth conditional control, we build the metrics in two ways:
+(1) Model-based metric.
+The consistency of depth maps estimated by large pre-trained models such as Midas (Ranftl et al.
+2020) and Zoe-Depth (Bhat et al.
+2023) can reflect the depth priors of the images.
+Following ControlNet++ (Li et al.
+2025), we build the M-MSE and Z-MSE metrics to evaluate the depth consistency between the input condition and the generated one.
+If Midas model (Ranftl et al.
+2020) is used, $f_{\text{cond}}=f_{\text{midas}}$.
+If Zoe-Depth (Bhat et al.
+2023) is used, $f_{\text{cond}}=f_{\text{zoe-depth}}$.
+The input conditional map $C_{\text{cond}}$ at viewpoint $r$ is $C_{d}$.
+The corresponding ground truth image is noted as $I_{r}^{\text{gt}}$.
+M-MSE This metric firstly extracts the relative depth maps from the images via pre-trained Midas model $f_{\text{midas}}$.
+Then, the MSE difference between the estimated depth maps is calculated as follows:
+$\text{M-MSE}(P,T_{r},I_{r}^{\text{gt}})=\text{MSE}(f_{\text{midas}}(f_{\text{render}}^{I}(P,T_{r})),f_{\text{midas}}(I_{r}^{\text{gt}}))$ (24)
+Z-MSE This metric firstly extracts the relative depth maps from the images via pre-trained Zoe-Depth model $f_{\text{zoe-depth}}$.
+Then, the MSE difference between the estimated depth maps is computed as follows:
+$\text{Z-MSE}(P,T_{r},I_{r}^{\text{gt}})=\text{MSE}(f_{\text{zoe-depth}}(f_{\text{render}}^{I}(P,T_{r})),f_{\text{zoe-depth}}(I_{r}^{\text{gt}}))$ (25)
+(2) 3D-based metric.
+Since the model-based metrics of M-MSE and Z-MSE only capture the depth consistencies from rendered 2D images, it is necessary to evaluate the consistency between the generated 3D representation and the 3D conditional control (depth).
+In consequence, the 3D-based metrics of R-MSE measures the MSE difference between the rendered depth map of the generated 3D content $P$ at viewpoint $r$ and the input depth condition map $C_{d}$.
+R-MSE This metric compares the difference between the rendered depth map and the input depth condition map as follows:
+$\text{R-MSE}(P,T_{r},C_{d})=\text{MSE}(f_{\text{norm}}(f_{\text{render}}^{D}(P,T_{r})),C_{d})$ (26)
+where $f_{\text{norm}}$ is the normalization function.
+Since the rendered depth map and the input depth condition map have different scales, it is necessary to align the scale of these depth maps.
+Assume that the input depth condition map is already normalized to a range of 0 to 1, we can also normalize the rendered depth map to the same range of 0 to 1 via $f_{\text{norm}}$.
+Controllability on Normal
+To evaluate the controllability of normal conditional control, we build the metrics in two ways:
+(1) Model-based metric.
+The consistency of normal maps estimated by pre-trained models like Normal-BAE (Bae, Budvytis, and Cipolla 2021) can reflect the normal priors of the images.
+Following ControlNet++ (Li et al.
+2025), we build NB-MSE metric to evaluate the normal consistency between the input normal condition and the one extracted from the generated result.
+The input normal condition is noted as $C_{n}$, and the ground truth image at viewpoint $r$ is noted as $I_{r}^{\text{gt}}$.
+We utilize the pre-trained model of Normal-BAE provided by ControlNet (Zhang, Rao, and Agrawala 2023), the condition extractor is: $f_{\text{cond}}=f_{\text{nb-mae}}$.
+NB-MSE This metric captures the similarity of normal priors captured by pre-trained model of Normal-BAE.
+$\text{NB-MSE}(P,T_{r},I_{r}^{\text{gt}})=\text{MSE}(f_{\text{nb-mae}}(f_{\text{render}}^{I}(P,T_{r})),f_{\text{nb-mae}}(I_{r}^{\text{gt}}))$ (27)
+(2) 3D-based metric.
+Besides the model-based metric, we further evaluate the depth-normal consistency between the directly rendered depth map and the input normal condition.
+It can capture the consistency in 3D space directly.
+Specifically, we utilize the metric of DN-CON to measure the depth-normal consistency based on the orthogonality between normal and local surface tangent (Yang et al.
+2018).
+DN-CON The rendered depth is converted to a normalized normal maps via a differentiable module (Huang et al. 2021), and be compared with the input normal condition as follows:
+$\text{DN-CON}(P,T_{r},C_{n})=\text{MSE}(f_{\text{d2n}}(f_{\text{render}}^{D}(P,T_{r})),C_{n})$ (28)
+Evaluation Metric on Generation Quality
+In the main manuscript, we also provide the quantitative comparison with other methods. Here we introduce the details of related metrics:
+FID In analogy with LATTE3D (Xie et al.
+2024), we compute the Fréchet Inception Distance (FID) (Heusel et al.
+2017) between the rendered images of the generated 3D object and the collected ground truth multi-view images in the evaluation benchmarks.
+This metric can reflect how well the generation results algin with the ground truth multi-view images in visual quality.
+CLIP-I Following MVControl (Li et al.
+2024), we also compute the CLIP score of image features extracted from the renderings of the generated 3D content and the collected ground truth images, noted as CLIP-I.
+This metric aims to measure the similarity between the rendering results of generated 3D objects and the ground truth images.
+CLIP-T Following MVControl (Li et al.
+2024), we also measure the CLIP score between image features extracted from the renderings and the text features extracted from the given text prompt.
+This metric aims to reveal the similarity between the generated 3D contents and the given text descriptions.
+View Settings The view setting of evaluation protocol in MVControl (Li et al.
+2024) only compute the metrics between the generated multi-view images and the ground turth image on the reference view.
+However, merely evaluating the performance on singe one view is not comprehensive for evaluating the 3D generated contents.
+In consequence, we utilize the following settings to build more comprehensive metrics: (1) All Views: All views are taken into account when calculating the FID, CLIP-I, and CLIP-T; (2) Front-K Views: Given the provision of only one reference view, the views on the back side may lack crucial cues for precise prediction, potentially leading to unreliable results in multi-view scenarios.
+Therefore, incorporating an additional evaluation of the views in front of the reference view is necessary.
+Consequently, we select the K=4 views closest to the given reference view for further metric computation.
+Baselines for Comparisons
+For comparison, we select a series of state-of-the-art methods as baselines in the controllable 3D generation task, including:
+  1. 1.
+    Optimization-based methods: GSGEN (Chen et al.
+2024), GaussianDreamer (Yi et al.
+2024), and DreamGaussian (Tang et al.
+2023a).
+These optimization-based methods iteratively distill the prior of large pre-trained generative diffusion models to generate the corresponding 3D contents.
+  2. 2.
+    Feed-forward-based methods: VolumeDiffusion (Tang et al.
+2023b), 3DTopia (Hong et al.
+2024).
+These models directly build a diffusion model on 3D representation like Triplane NeRF (Mildenhall et al.
+2021), and generate 3D contents via the guidance of the input condition on the diffusion model.
+  3. 3.
+    Controllable 3D generation methods: MVControl (Li et al.
+2024), and ControLRM (Xu et al.
+2024a).
+These models are state-of-the-art models built for controllable 3D generation.
+They accept conditional maps and text prompts as input, and generate corresponding 3D contents via the pre-trained model.
+Below we provide the details needed to reproduce the baseline methods shown in the paper.
+GSGEN We use the  [^1] https://github.com/gsgen3d/gsgenofficial code provided by GSGEN (Chen et al.
+2024).
+The text prompts in the evaluation benchmarks are used as the input to generate the 3D contents.
+The default hyper-parameters of GSGEN are used.
+GaussianDreamer We use the  [^2] https://github.com/hustvl/GaussianDreamerofficial code provided by the GaussianDreamer (Yi et al.
+2024).
+The text prompts in the evaluation benchmarks are set as input to generate the 3D assets.
+The default hyper-parameters in their official implementation are used.
+DreamGaussians We use the  [^3] https://github.com/dreamgaussian/dreamgaussianofficial code of DreamGaussian (Tang et al.
+2023a).
+The text prompts in the evaluation benchmarks are used as the input for generating the 3D contents.
+The default hyper-parameters of their official implementation are directly used for experiments.
+VolumeDiffusion We use the  [^4] https://github.com/PeterZs/VolumeDiffusionofficial code of VolumeDiffusion (Tang et al.
+2023b).
+The text prompts of the evaluation datasets are used as input.
+We use the  [^5] https://github.com/PeterZs/VolumeDiffusion?tab=readme-ov-file#inferencepre-trained model of VolumeDiffusion to generate the 3D contents.
+Figure 7: Illustration of Pseudo Cycle Back-Propagation for implementing the cycle consistency with less GPU memory costs.
+[…]
+FID CLIP-I CLIP-T FID CLIP-I CLIP-T FID CLIP-I CLIP-T
+GSGEN (Chen et al. 2024) 340.39 0.762 0.298 344.61 0.740 0.289 366.47 0.669 0.259
+GaussianDreamer (Yi et al. 2024) 268.95 0.803 0.309 278.7 0.810 0.300 225.38 0.787 0.277
+DreamGaussian (Tang et al. 2023a) 351.87 0.761 0.279 359.65 0.760 0.279 392.95 0.723 0.247
+VolumeDiffusion (Tang et al. 2023b) 327.76 0.725 0.241 299.61 0.715 0.259 350.46 0.679 0.242
+3DTopia (Hong et al. 2024) 289.02 0.749 0.280 331.39 0.727 0.283 231.55 0.751 0.272
+MVControl (Li et al. 2024) 251.71 0.811 0.291 278.08 0.816 0.288 217.97 0.802 0.291
+ControLRM-T (Xu et al. 2024a) 166.03 0.879 0.292 260.75 0.846 0.289 202.14 0.827 0.282
+ControLRM-D (Xu et al. 2024a) 163.25 0.887 0.300 171.13 0.838 0.302 181.84 0.836 0.292
+CyC3D-T (Ours) 170.84 0.880 0.292 255.30 0.840 0.295 196.55 0.836 0.291
+CyC3D-D (Ours) 171.21 0.881 0.297 168.71 0.846 0.303 180.34 0.843 0.294
+Table 7: Generation quality (FID) and CLIP score (CLIP-I, CLIP-T) comparison under All View setting with other state-of-the-art controllable 3D generation method on GOBJ, GSO, and ABO benchmarks.
+$\uparrow$ means higher result is better, while $\downarrow$ means lower is better.
+[…]
+FID CLIP-I CLIP-T FID CLIP-I CLIP-T FID CLIP-I CLIP-T
+GSGEN (Chen et al. 2024) 357.13 0.781 0.310 360.57 0.759 0.300 376.56 0.691 0.272
+GaussianDreamer (Yi et al. 2024) 282.59 0.823 0.321 287.20 0.829 0.311 226.56 0.831 0.306
+DreamGaussian (Tang et al. 2023a) 368.76 0.783 0.293 373.53 0.784 0.290 406.35 0.750 0.290
+VolumeDiffusion (Tang et al. 2023b) 348.39 0.752 0.257 316.35 0.742 0.273 372.49 0.715 0.262
+3DTopia (Hong et al. 2024) 329.29 0.808 0.312 369.27 0.799 0.311 259.88 0.844 0.312
+MVControl (Li et al. 2024) 280.40 0.856 0.318 301.31 0.870 0.312 236.81 0.868 0.316
+ControLRM-T (Xu et al. 2024a) 144.02 0.932 0.323 251.57 0.912 0.316 160.12 0.915 0.320
+ControLRM-D (Xu et al. 2024a) 148.76 0.935 0.330 247.06 0.908 0.322 152.37 0.918 0.324
+CyC3D-T (Ours) 155.08 0.930 0.322 247.03 0.913 0.320 159.03 0.904 0.322
+CyC3D-D (Ours) 155.82 0.931 0.327 215.98 0.915 0.324 154.85 0.908 0.324
+Table 8: Generation quality (FID) and CLIP score (CLIP-I, CLIP-T) comparison under Front-K View setting with other state-of-the-art controllable 3D generation method on GOBJ, GSO, and ABO benchmarks.
+$\uparrow$ means higher result is better, while $\downarrow$ means lower is better.
+3DTopia We use the  [^6] https://github.com/3DTopia/3DTopiaofficial code of 3DTopia (Hong et al.
+2024).
+The text prompts in the evaluation datasets are used as input.
+We use their open-sourced  [^7] https://huggingface.co/hongfz16/3DTopiapre-trained checkpoints for experiments.
+The default parameters in the official code are used.
+MVControl We use the  [^8] https://github.com/WU-CVGL/MVControlofficial implementation of the state-of-the-art controllable 3D generation model, MVControl (Li et al.
+2024).
+The open-source pre-trained checkpoints of different conditional controls are used for experiments, including:  [^9] https://huggingface.co/lzq49/mvcontrol-4v-cannyEdge (Canny),  [^10] https://huggingface.co/lzq49/mvcontrol-4v-depthDepth,  [^11] https://huggingface.co/lzq49/mvcontrol-4v-normalNormal,  [^12] https://huggingface.co/lzq49/mvcontrol-4v-scribbleSketch.
+The condition map on the reference view and the corresponding text prompt are fed to the model as input.
+The same parameters in the official code are used for a fair comparison.
+ControLRM We use the  [^13] https://toughstonex.github.io/controlrm.github.io/official implementation of ControLRM (Xu et al.
+2024a).
+the open-source pre-trained checkpoints of different conditional controls are used for experiments, including:  [^14] https://toughstonex.github.io/controlrm.github.io/Edge (Canny),  [^15] https://toughstonex.github.io/controlrm.github.io/Depth,  [^16] https://toughstonex.github.io/controlrm.github.io/Normal,  [^17] https://toughstonex.github.io/controlrm.github.io/Sketch.
+The condition map on the reference view and the corresponding text prompt are fed to the model as input.
+The same parameters in the official code are used for a fair comparison.
+Implementation Details
+Backbone Network In analogy with (Xu et al.
+2024a), we initialize our network with the weights from pretrained OpenLRM-base (Hong et al.
+2023).
+The image-conditioned transformer is removed and replaced with our customized conditional backbone network taking visual conditions (edge/sketch/depth/normal) as input.
+The output feature is fed to the activated cross-attention layers in the triplane transformer of OpenLRM, while the remaining parameters are kept frozen.
+ControLRM (Xu et al.
+2024a) has 2 variations with different conditional backbones: ControLRM-T with transformer backbone, and ControLRM-D with diffusion backbone.
+In analogy, we inherit the same conditional backbone of ControLRM to our CyC3D, and note them as CyC3D-T/-D rescpectively.
+Training Details We utilize 32 Nvidia V100-32G GPUs for training CyC3D-T and CyC3D-D.
+For CyC3D-T, the training might cost 2 days.
+For CyC3D-D, the training might cost 3 days.
+We utilize the AdamW optimizer with a learning rate of 4e-4.
+The resolution of the condition map (edge/sketch/depth/normal) is set to $336\times 336$.
+To save GPU memories, the resolution of the rendered image is set to $256\times 256$ in default.
+The training code is heavily based on the open-source project  [^18] https://github.com/3DTopia/OpenLRMOpenLRM, and will be release in the future.
+Pseudo Cycle Back-Propagation In the implementation, a non-negligible issue is the huge memory cost during training.
+The volume rendering has to be run twice in the cycle consistency, hence even the V100-32G GPU might suffer from the problem of GPU memory overflow.
+To handle this problem, we propose Pseudo Cycle Back-Propagation to trade space with time, as shown in Fig.
+7.
+The orange arrows in the figure denotes the back-ward gradients in the computation graph.
+Instead of calculating the whole computation graph in one round, we separate it into two rounds as the figure shows.
+In the first round, we back-propagate the gradients to the intermediate output (Generated conditions at randomly sampled views).
+The gradients on the intermediate output are cached and saved as tensors.
+In the second round, we back-propagate these cached gradients to the corresponding nodes in the computation graph, and finish the remaining process of back-propagation.
+In summary, we replace the original computation pipeline (one forward-propagation, one backward-propagation) with our modified computation pipeline (one forward-propagation, two backward-propagation).
+It can effectively reduce the GPU memory cost during the training phase.
+Additional Experimental Results
+Time Efficiency
+Figure 8: Time efficiency comparison with other methods in controllable 3D generation.
+In Fig.
+8, we compare our CyC3D with other methods in time efficiency.
+The inference time of the model on a single V100-32G GPU is reported in the figure.
+Since our model utilizes the same backbone as ControLRM (Xu et al.
+2024a), our proposed method will not add extra inference budget to the backbone network meantime improving the controllability quality.
+Additional Quantitative Results on Generation Quality
+The controllable 3D generation model should not only improve the controllability, but also maintain a good generation quality.
+Usually, there exists a trade-off between these two aspects depending on which property one prefers.
+In consequence, we also provide the quantitative comparison of the generation quality here.
+As discussed in Sec.
+Evaluation Metric on Generation Quality, we compare different method on FID/CLIP-I/CLIP-T under two different view settings (All View/Front-K View).
+The quantitative results under All View setting are presented in Tab.
+7.
+The quantitative results under Front-K Views setting are shown in Tab.
+8.
+Our model can maintain competitive and even better generation quality under zero-shot generalization evaluation, i.e.
+$\rightarrowGSO, GOBJ\rightarrowABO$.
+But the generation quality on the same dataset might be slightly worse than the existing SOTA.
+This implies that our model can generalize to unseen domains well by sacrificing the over-fitting effect to the training domain.
+Comparison under Different Conditional Controls
+Metrics Methods Edge Depth Normal Sketch Means
+Control MVControl (Li et al. 2024) 0.0952 0.0160 0.0315 0.0370 0.0449
+ControLRM-T (Xu et al. 2024a) 0.0607 0.0599 0.0359 0.0202 0.0442
+ControLRM-D (Xu et al. 2024a) 0.0600 0.0650 0.0357 0.0197 0.0451
+CyC3D-T (Ours) 0.0342 0.0065 0.0062 0.0159 0.0157
+CyC3D-D (Ours) 0.0347 0.0060 0.0060 0.0156 0.0156
+FID MVControl (Li et al. 2024) 311.91 256.08 255.68 288.64 278.08
+ControLRM-T (Xu et al. 2024a) 253.59 262.95 269.62 256.84 260.75
+ControLRM-D (Xu et al. 2024a) 165.38 174.29 174.98 169.88 171.13
+CyC3D-T (Ours) 258.16 248.84 250.87 263.32 255.30
+CyC3D-D (Ours) 168.11 168.49 167.78 170.46 168.71
+CLIP-I MVControl (Li et al. 2024) 0.762 0.841 0.851 0.811 0.816
+ControLRM-T (Xu et al. 2024a) 0.855 0.842 0.835 0.852 0.846
+ControLRM-D (Xu et al. 2024a) 0.854 0.826 0.820 0.850 0.838
+CyC3D-T (Ours) 0.847 0.828 0.847 0.839 0.840
+CyC3D-D (Ours) 0.848 0.846 0.843 0.846 0.846
+CLIP-T MVControl (Li et al. 2024) 0.263 0.298 0.299 0.291 0.288
+ControLRM-T (Xu et al. 2024a) 0.293 0.288 0.284 0.291 0.289
+ControLRM-D (Xu et al. 2024a) 0.304 0.301 0.300 0.303 0.302
+CyC3D-T (Ours) 0.296 0.296 0.295 0.292 0.295
+CyC3D-D (Ours) 0.303 0.304 0.303 0.302 0.303
+[…]
+[…]
+Additional Ablation Results
+Figure 9: Failure Cases.
+Figure 10: More visualization results.
+Ablation of Different Backbones To find out the performance when different backbones are adopted, we can respectively compare CyC3D-T/-D with ControLRM-T/-D in Tab.
+9.
+Our CyC3D adopts the same network architectures as ControLRM under 2 different network backbones (transformer/diffusion), noted as -T/-D.
+From the experimental results, we can find that both CyC3D-T and CyC3D-D achieve better controllability scores compared with ControLRM-T and ControLRM-D.
+Furthermore, the generation quality (FID, CLIP-I, CLIP-T) of our CyC3D-T and CyC3D-D is competitive and even better than ControLRM-T and ControLRM-D.
+It demonstrates the effectiveness of our proposed method on different backbone networks, including transformer and diffusion backbones.
+Loss Controllability Generation Quality
+PSNR SSIM MSE FID CLIP-I CLIP-T
+w/o $L_{\text{cond}}$ 12.87 0.841 0.0570 165.38 0.854 0.304
+w $L_{\text{cond}}$ 15.09 0.885 0.0347 168.71 0.846 0.303
+[…]
+Loss Controllability Generation Quality
+PSNR SSIM MSE FID CLIP-I CLIP-T
+w/o $L_{\text{cond}}$ 17.52 0.888 0.0187 169.88 0.850 0.303
+w $L_{\text{cond}}$ 18.38 0.890 0.0156 170.46 0.846 0.302
+[…]
+Loss Controllability Generation Quality
+M-MSE Z-MSE R-MSE FID CLIP-I CLIP-T
+w/o $L_{\text{cond}}$ 0.0383 0.0546 0.0650 174.29 0.826 0.301
+w $L_{\text{cond}}$ 0.0192 0.0241 0.0321 172.10 0.832 0.302
+w $L_{\text{cond}}+L_{\text{cond-d}}$ 0.0051 0.0041 0.0060 168.49 0.846 0.304
+[…]
+Loss Controllability Generation Quality
+NB-MSE DN-CON FID CLIP-I CLIP-T
+w/o $L_{\text{cond}}$ 0.0034 0.0205 174.98 0.820 0.300
+w $L_{\text{cond}}$ 0.0023 0.0187 172.18 0.831 0.302
+w $L_{\text{cond}}+L_{\text{cond-n}}$ 0.0011 0.0056 167.78 0.843 0.303
+[…]
+[…]
+[…]
+[…]
+[…]
+Discussion
+Improvement on Depth/Normal vs Edge/Sketch From the quantitative and qualitative results, we can observe that: the improvement of our proposed CyC3D on depth/normal condition control is much higher than the ones on edge/sketch conditional control.
+We summarize two important issues as follows:
+(1) Alignment of correct 3D condition maps with the ground truth 3D object.
+In existing methods (like MVControl (Li et al.
+2024)), the depth and normal condition maps are usually extracted from images with pre-trained models.
+However, these models tend to predict over-smoothed depth or normal maps.
+Furthermore, these generated depth or normal maps have different scales with the real 3D object, leading to incorrect 3D shape in these generated condition maps.
+Since these incorrect depth or normal maps are set as input, the training phase will enforce the generated 3D shapes to align with these incorrect 3D conditions.
+Consequently, this problem may significantly degrade the final performance of controllable 3D generation models.
+Different from existing methods, our CyC3D can effectively align the 3D representation and the input 3D conditons like depth and normal.
+It can not only mitigate the difference in scale, but also avoid the discordance between the 3D condtion and the 3D object.
+Consequently, our CyC3D can improve the controllability significantly meantime providing competitive and even better generation quality.
+(2) 2D condition maps like edge/sketch may have ambiguities when mapping from 2D to 3D.
+During the preparation edge or sketch condition maps in the dataset, the extracted edge or sketch conditions are usually noisy and sensitive to hyper-parameters or disturbances in images like light conditions.
+This leads to an unexpected fact that the noisy condition maps might contain ambiguous information.
+For example, some reflection of lights might leads to fake edges or sketchs in the condition maps, some noise in the images might create dots-like edges or sketchs in the condition maps.
+In consequence, the natural disturbance in edge/sketch maps might involve extra noise during training, thus leading to degradation of generation quality slightly.
+But the controllability is not affected by this issue, because it only concentrates on the cycle consistency between input condition and generated results.
+Hence, the controllability is improved significantly normally.
+Difference with Cycle3D Note that Cycle3D (Tang et al.
+2025b) also applies the idea of cycle consistency in image-to-3D generation tasks, achieving superior performance.
+The differences between our work and Cycle3D are summarized as follows: (1) Cycle3D builds the cycle consistency on the pipeline of: diffusion model $\rightarrow$ reconstruction model $\rightarrow$ diffusion model –¿ … It is an inference pipeline for generating corresponding 3D contents from given image inputs.
+However, our work is totally different.
+Our CyC3D applies the cycle consistency constraint of conditional inputs during the training phase.
+Whereas our CyC3D directly generates the results with one feed-forward model during inference, rather than propagating through multiple rounds of diffusion and reconstruction models in Cycle3D.
+(2) Our CyC3D handles a different task compared with Cycle3D.
+We aim to handle the controllable 3D generation task with conditional inputs like edge, depth, normal or sketch maps.
+Whereas Cycle3D only handles the image-to-3D generation problem.
+Failure Cases In Fig.
+9, we show some failure cases of our model.
+The first column is the text prompt, and the second column is the input condition map.
+The remaining columns visualize the rendered images and depths at different viewpoints.
+(1) As the the first row shows, due to the strong regularization on the controllability of our CyC3D, the blurry and noisy edge condition in the blue dotted box might make the model confusing to generate a reasonable result.
+(2) As the second row shows, the generated result seems to be reasonable on the same reference view, but fails at other viewpoints.
+The blue dotted box highlights the failures of the generated 3D object.
+Limitation Based on aforementioned analysis on failure cases in Fig.
+9, we can summarize the limitation of our proposed as follows: (1) Our model requires accurate conditional control, otherwise the blurry conditions might lead to unsatisfying generated results.
+Because the strong controllability ability of CyC3D might reversely be confused by inaccurate or blurry condition maps.
+A future direction is how to mitigate these disturbances in condition maps.
+(2) The bottleneck of our model depends on the bottleneck of the pre-trained backbone (i.e.
+OpenLRM).
+Since we directly inherits the parameters of triplane transformers in OpenLRM-base, the bottleneck of the backbone model (OpenLRM-base) limits our model as well.
+The performance of our model depends on the performance of the backbone model.
+Replacing the backbone with a large and powerful model can increase the performance significantly.